Prove that a Hilbert space is seperableiff every ortho normal set in

H is countable.


1
Expert's answer
2021-12-02T12:17:35-0500

1. If H is separable, every orthonormal set in H is countable

2. If H contains an orthonormal sequence which is total in H, then H is separable.

Proof:

1.

Let B be a countable dense subset of H. Let M be an uncountable orthonormal set. Then the distance between any two elements x, y ∈ M is √ 2 since

"||x-y||^2=||x||^2+||y||^2=2"

For each x ∈ M, we define Nx as the ball centered at x with radius √ 2/3. Then the Nx are disjoint. Since B is dense, for each x, there is bx ∈ B such that bx ∈ Nx. Since Nx are disjoint, the collection {bx} is an uncountable subset of B. This is a contradiction.


2.

Let A be the set of all linear combinations

"\u03b3 _1^{(n) } e_1 + . . . +\u03b3 _n^{(n) } e_n , n = 1, 2,..."

where the coefficients γ are complex rational (γ = a + ib, both a and b rational) or just rational (if the underlying field was real). Think of A as the rational-span of (ek).

Then A is the countable union of countable sets. We claim it is dense in H. Fix any x ∈ H and any

"\varepsilon" > 0. Then since the sequence (en) is total in H, there is some point

"y = \sum \langle e_k,x\rangle e_k"

such that "||k_x − y_k|| < \varepsilon/2" , and then se use the triangle inequality and the density of the rationals to see that there is a nearby point v such that "||x − v ||<\varepsilon"